1)     It has been determined that 4% of drivers checked at a road stop show traces of alcohol and 11% of drivers checked do not wear seat belts. In addition, it has been observed that the two infractions are independent of one another. 

A)  Calculate the probability that none of the drivers show traces of alcohol.


1
Expert's answer
2021-04-20T07:56:32-0400

Probability that driver shows traces of alcohol P(A)=4%=0.04

Probability that driver do not wear seat belts P(B)=11%=0.11


Probability that none of the driver shows traces of Alcohol -

"=1-P(A\cap B)
\\
 =1-P(A).P(B)\\=1-(0.04)(0.11)\\=1-0.0044=0.9956"